# Instructions for Authors: Title in Title Case

### v1

@@ B.4.3 Results of IBCB’s Parameter Test on Synthetic Dataset
B-IRL 0.588±0.015 0.571±0.037 0.572±0.034 0.561±0.034
BaseSVC 0.577±0.017 0.559±0.042 0.559±0.037 0.547±0.037
B-ICB-S 0.581±0.006 0.553±0.024 0.560±0.025 0.552±0.025
IBCB-$\alpha$-1 0.598±0.017 0.580±0.042 0.580±0.039 0.569±0.038
IBCB-$\alpha$-0.6 0.598±0.017 0.580±0.042 0.580±0.039 0.569±0.038
IBCB-$\alpha$-0.8 0.598±0.017 0.580±0.042 0.580±0.039 0.569±0.038
IBCB-$\alpha$-1.2 0.598±0.017 0.580±0.042 0.580±0.039 0.569±0.038
IBCB-$\alpha$-1.5 0.599±0.016 0.580±0.042 0.580±0.040 0.569±0.038
IBCB-$\alpha$-2 0.599±0.016 0.580±0.042 0.580±0.040 0.569±0.038
‘IBCB-$\alpha$-k’ denotes IBCB’s hyper parameter (in Eq.(10)) is k. As shown in Table 13 and Table 14, IBCB’s hyper parameter $\alpha$ will influence several indicators, especially the OL-Fitness and train time, due to the OSQP tolerance level will be relaxed if $\alpha$ is above 1.5. So for IBCB of large $\alpha$ with more relaxed tolerance, $i.e.,$ lower accuracy, training will be faster. Meanwhile, the performance of expert policy’s parameter estimation is similar according to BT-Fitness and BT-AR. By examining the parameters learned by IBCB with different $\alpha$ values, we found that after normalization (by dividing each parameter’s $l_{2}$-norm), the sum of absolute differences in every dimension of the parameters from each pair of two distinct parameters is tightly small. This suggests that one of these parameters can be a result of scaling up or down from another. During BT-data testing, items are recommended based on the parameter under BCB setting without exploration. Therefore, scaling up or down does not influence the recommended item since $\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\hat{\bm{\theta}},\bm{s}_{I}\rangle$ is only determined by the item’s content $\bm{s}_{I}$ when $\hat{\bm{\theta}}$ is fixed with a scaling factor. On the synthetic dataset, the item’s content $\bm{s}_{I}$ is very easy to separate (interval difference of each dimension of $\bm{s}_{I}$ is very obvious through the setting in Appendix B.1), so two distinct parameters with very small sum of absolute differences may share almost the same BT-Fitness and BT-AR.
B.4.4 Results of Ablation Study on ML-100K Dataset
@@
Table 15: Reward & policy parameters’ estimation on ML-100K dataset (using first half of expert’s behavioral evolution history logs). Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1. BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison.
@@
IBCB (Ours) 0.843±0.019 0.821±0.027 0.191±0.006
Table 16: IBCB’s parameter test on ML-100K dataset. Experts do not need to train since they produce train logs (behavioral evolution history logs) for baselines and IBCB.
@@
IBCB-$\alpha$-1 0.870±0.014 0.884±0.021 0.192±0.006 15.023±3.854
IBCB-$\alpha$-0.6 0.829±0.019 0.884±0.021 0.192±0.006 14.942±3.807
IBCB-$\alpha$-0.8 0.854±0.017 0.884±0.021 0.192±0.006 14.902±3.813
IBCB-$\alpha$-1.2 0.880±0.015 0.880±0.024 0.192±0.006 62.138±59.186
IBCB-$\alpha$-1.5 0.883±0.014 0.876±0.022 0.192±0.006 5.647±0.859
IBCB-$\alpha$-2 0.888±0.010 0.879±0.023 0.192±0.006 6.722±1.641
For ML-100K dataset, Table 15 and Table 16 reports the result of different ablation studies. Results are similar to those on synthetic dataset in most indicators: IBCB can learn from novice expert’s evolution history with better performance compared with other baselines. Note that mean and variance of train time for IBCB-$\alpha$-1.2 is very large. From the OSQP train record, we analysed that this happened due to OSQP failed to convergent for some of the OL-data train logs, so for large $\alpha$, it is recommended to use more relaxed tolerance level in OSQP setting. Note that B-ICB-S’s performance significantly decline when using first half of the train logs in Table 15, this is because when in first half of the real-world train logs, novice expert is still exploring with large steps in a batched manner rather than updating after each step or staying stable with little exploration, which B-ICB-S assumed.
B.4.5 Results of Extra Ablation Study Experiments on ML-100K Dataset

### v2

@@ 5.6.2 Limited Training logs of Experts on Synthetic Dataset
B-IRL 0.545±0.020 0.522±0.047 0.522±0.046 0.510±0.046
BaseSVC 0.520±0.023 0.497±0.053 0.497±0.052 0.482±0.051
B-ICB-S 0.549±0.019 0.538±0.024 0.540±0.030 0.529±0.033
IBCB (Ours) 0.610±0.015 0.608±0.025 0.597±0.036 0.588±0.030
From the results in Table VII and Table VIII we find that, when all IL baselines and IBCB only have the access to the first half of expert’s behavioral evolution history logs, i.e., novice expert’s evolution history, IBCB can have better performance, compared with the degree of decline in all algorithms from Table I and Table II.
5.6.3 Extra Fairness-aware Expert Experiments on Synthetic Dataset
TABLE IX: Cumulative Fairness Regret (CFR) comparison (Upper Subtable) and Train time comparison (Lower Subtable) on synthetic dataset. Expert selects action from top-$\bm{3}$ actions with proportion of its learned probability distribution through all actions. Expert is defined as the fairest policy from the CFR metric, so its CFR is always 0. Fairness-aware experts (SBUCB & SBTS) do not need to train since they produce the train logs (behavioral evolution history logs) for baselines and IBCB.
Algorithm Cumulative Fairness Regret Algorithm Cumulative Fairness Regret
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 0 0 0 0 SBTS 0 0 0 0
B-IRL 2713.9±230.2 2882.3±356.2 2841.7±356.4 2982.029±370.4 B-IRL 2704.2±300.9 2433.9±285.2 2498.8±248.0 2488.4±226.6
BaseSVC 2465.2±257.4 2653.5±412.5 2608.6±447.9 2758.4±421.7 BaseSVC 2395.3±346.8 2131.7±376.1 2166.1±291.1 2170.2±299.2
B-ICB-S 2637.6±168.5 2747.3±260.1 2805.3±398.6 2631.2±485.6 B-ICB-S 2856.4±439.8 3281.5±162.1 3136.9±187.3 2980.8±350.3
IBCB (ours) 2232.6±148.6 2071.0±264.4 2161.9±124.4 2192.0±110.0 IBCB (ours) 1596.7±172.3 1704.7±158.6 1584.4±124.1 1705.5±176.8
Algorithm Train Time (sec.) Algorithm Train Time (sec.)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB / / / / SBTS / / / /
B-IRL 40.758±4.040 35.796±0.352 38.59±4.594 41.948±7.858 B-IRL 35.664±0.208 37.07±2.135 39.162±6.970 35.662±0.081
BaseSVC 234.56±13.51 228.80±19.52 231.25±31.15 249.73±31.44 BaseSVC 209.01±10.75 221.06±18.02 215.58±18.12 230.40±25.84
B-ICB-S 85.977±1.975 87.631±4.364 86.164±1.884 84.881±0.493 B-ICB-S 86.726±1.870 87.447±1.977 86.796±1.808 85.974±1.826
IBCB (ours) 1.169±0.225 1.102±0.084 1.075±0.048 1.055±0.031 IBCB (ours) 0.798±0.020 0.786±0.012 0.791±0.013 0.849±0.116
TABLE X: Fairness-aware expert policy (SBUCB) parameters’ estimation (Upper Subtable) and Fairness-aware expert policy (SBTS) parameters’ estimation (Lower Subtable) on synthetic dataset. Expert selects action from top-$\bm{3}$ actions with proportion of its learned probability distribution through all actions. Expert policy produces same actions compared with itself, so Batch Test Log Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness Algorithm Batch Test Average Reward
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 1 1 1 1 SBUCB 0.538±0.012 0.529±0.030 0.528±0.031 0.516±0.031
B-IRL 0.307±0.005 0.297±0.008 0.300±0.008 0.296±0.007 B-IRL 0.486±0.018 0.472±0.039 0.472±0.041 0.457±0.042
BaseSVC 0.306±0.006 0.302±0.01 0.305±0.009 0.303±0.009 BaseSVC 0.481±0.021 0.467±0.041 0.467±0.044 0.451±0.043
B-ICB-S 0.300±0.006 0.296±0.006 0.295±0.008 0.296±0.011 B-ICB-S 0.500±0.004 0.484±0.025 0.482±0.029 0.475±0.026
IBCB (ours) 0.307±0.004 0.310±0.006 0.309±0.004 0.309±0.004 IBCB (ours) 0.558±0.010 0.543±0.029 0.545±0.027 0.535±0.029
Algorithm Batch Test Log Fitness Algorithm Batch Test Average Reward
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBTS 1 1 1 1 SBTS 0.548±0.028 0.566±0.030 0.560±0.025 0.556±0.024
B-IRL 0.302±0.006 0.309±0.003 0.305±0.002 0.307±0.002 B-IRL 0.495±0.039 0.521±0.037 0.512±0.030 0.508±0.028
BaseSVC 0.308±0.007 0.311±0.005 0.312±0.006 0.312±0.004 BaseSVC 0.492±0.039 0.518±0.039 0.510±0.032 0.507±0.031
B-ICB-S 0.292±0.011 0.281±0.008 0.287±0.004 0.292±0.009 B-ICB-S 0.501±0.014 0.493±0.021 0.496±0.016 0.496±0.017
IBCB (ours) 0.321±0.004 0.319±0.004 0.322±0.003 0.318±0.004 IBCB (ours) 0.547±0.031 0.568±0.031 0.559±0.026 0.559±0.022
TABLE XI: Cumulative Fairness Regret (CFR) comparison (Upper Subtable) and Train time comparison (Lower Subtable) on synthetic dataset. Expert selects action from top-$\bm{4}$ actions with proportion of its learned probability distribution through all actions. Expert is defined as the fairest policy from the CFR metric, so its CFR is always 0. Fairness-aware experts (SBUCB & SBTS) do not need to train since they produce the train logs (behavioral evolution history logs) for baselines and IBCB.
Algorithm Cumulative Fairness Regret Algorithm Cumulative Fairness Regret
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 0 0 0 0 SBTS 0 0 0 0
B-IRL 3063.9±258.1 3279.7±364.6 3214.4±381.0 3263.6±286.0 B-IRL 2935.1±407.0 2812.5±254.6 2790.2±260.6 2794.1±319.6
BaseSVC 2893.6±240.7 3086.1±401.7 3042.2±399.8 3121.7±302.5 BaseSVC 2752.0±379.1 2597.6±225.7 2617.2±273.2 2588.8±261.7
B-ICB-S 2998.0±153.3 3109.8±364.9 3123.4±287.4 3018.3±165.1 B-ICB-S 3295.3±357.9 3092.2±240.2 3369.5±59.5 3093.9±388.0
IBCB (ours) 3130.6±194.2 2794.3±282.2 2966.9±205.4 2882.1±238.9 IBCB (ours) 2162.4±129.4 2117.3±43.8 1918.0±128.6 1990.1±232.6
Algorithm Train Time (sec.) Algorithm Train Time (sec.)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB / / / / SBTS / / / /
B-IRL 36.867±0.571 43.7±6.185 40.665±6.652 37.241±0.991 B-IRL 36.697±0.373 41.568±4.424 39.308±5.713 36.270±0.706
BaseSVC 330.68±22.03 343.56±37.87 349.05±37.32 365.68±7.47 BaseSVC 295.92±23.73 327.44±37.10 314.77±19.73 309.99±14.98
B-ICB-S 86.362±1.885 84.884±0.443 86.48±0.594 86.488±1.541 B-ICB-S 88.003±2.241 86.189±2.22 86.318±1.266 87.87±1.607
IBCB (ours) 1.052±0.052 1.016±0.037 1.135±0.211 1.034±0.032 IBCB (ours) 0.950±0.175 0.737±0.013 0.838±0.148 0.820±0.112
TABLE XII: Fairness-aware expert policy (SBUCB) parameters’ estimation (Upper Subtable) and Fairness-aware expert policy (SBTS) parameters’ estimation (Lower Subtable) on synthetic dataset. Expert selects action from top-$\bm{4}$ actions with proportion of its learned probability distribution through all actions. Expert policy produces same actions compared with itself, so Batch Test Log Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness Algorithm Batch Test Average Reward
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 1 1 1 1 SBUCB 0.490±0.01 0.480±0.029 0.478±0.029 0.471±0.028
B-IRL 0.236±0.003 0.229±0.004 0.229±0.005 0.230±0.005 B-IRL 0.444±0.014 0.429±0.035 0.428±0.038 0.421±0.031
BaseSVC 0.235±0.003 0.232±0.007 0.232±0.005 0.233±0.006 BaseSVC 0.441±0.016 0.424±0.038 0.423±0.038 0.414±0.032
B-ICB-S 0.228±0.003 0.226±0.007 0.227±0.007 0.227±0.004 B-ICB-S 0.456±0.007 0.450±0.024 0.443±0.024 0.442±0.022
IBCB (ours) 0.227±0.003 0.231±0.007 0.230±0.003 0.230±0.004 IBCB (ours) 0.523±0.005 0.499±0.024 0.502±0.024 0.495±0.024
Algorithm Batch Test Log Fitness Algorithm Batch Test Average Reward
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBTS 1 1 1 1 SBTS 0.500±0.024 0.510±0.021 0.510±0.022 0.503±0.026
B-IRL 0.234±0.006 0.234±0.003 0.235±0.002 0.233±0.002 B-IRL 0.455±0.033 0.466±0.026 0.465±0.026 0.460±0.032
BaseSVC 0.235±0.004 0.238±0.005 0.235±0.003 0.235±0.004 BaseSVC 0.453±0.034 0.466±0.027 0.464±0.028 0.458±0.032
B-ICB-S 0.226±0.006 0.229±0.007 0.222±0.007 0.228±0.009 B-ICB-S 0.451±0.016 0.465±0.021 0.462±0.016 0.458±0.018
IBCB (ours) 0.239±0.004 0.241±0.002 0.242±0.004 0.241±0.003 IBCB (ours) 0.496±0.024 0.507±0.020 0.500±0.016 0.499±0.032
We also conducted fairness-aware expert experiments with different fairness settings on synthetic dataset. Specifically, in the original experiment, we set top-$k$ fairness with $k=2$ (Table IV–Table V). We also make experiments with $k=3,4$ to discover the impact of changing fairness parameter (Table IX-Table XII). It can be observed from the results that when $k$ increases, BC methods’ training time significantly increase, indicating that BC methods tend to be confused because observed actions made by the expert are drawn from more possible candidates, i.e. the top-$k$ set. When $k$ increased, the BT-Fitness performance of all models decreased significantly, and the CFR also increased rapidly, which is related to the more random action selection of experts. However, IBCB can still quickly learn expert parameters and get the best performance in this situation, which also shows that IBCB still has a very high generalization ability in complex scenarios of fairness.
5.6.4 Detailed Results of Contradictory Data Testing on Synthetic Dataset
TABLE XIII: Contradictory (duplicated) data (appeared at expert’s OL-data behavioral evolution history logs) on synthetic dataset testing. $\mathrm{dup}$ denotes the quantity of duplicated episodes in OL-data phase. This table shows the OL-Fitness results.
Algorithm Online Train Log Fitness
dup=0 dup=1 dup=2 dup=3 dup=4 dup=5 dup=6 dup=7 dup=8 dup=9
SBUCB 0.879±0.016 0.880±0.015 0.880±0.015 0.879±0.015 0.880±0.018 0.879±0.015 0.880±0.016 0.879±0.015 0.879±0.015 0.878±0.017
B-IRL 0.808±0.039 0.808±0.037 0.808±0.038 0.806±0.038 0.807±0.038 0.807±0.037 0.809±0.037 0.810±0.037 0.809±0.037 0.808±0.037
IBCB (Ours) 0.858±0.034 0.860±0.021 0.855±0.034 0.865±0.021 0.868±0.005 0.874±0.003 0.875±0.006 0.848±0.038 0.871±0.006 0.867±0.008
Algorithm Train Log Behaviors Fitness (BC)
dup=0 dup=1 dup=2 dup=3 dup=4 dup=5 dup=6 dup=7 dup=8 dup=9
BaseSVC 0.823±0.026 0.825±0.025 0.824±0.026 0.821±0.026 0.820±0.026 0.819±0.025 0.819±0.025 0.819±0.024 0.817±0.023 0.816±0.023
TABLE XIV: Contradictory (duplicated) data (appeared at expert’s OL-data behavioral evolution history logs) on synthetic dataset testing. $\mathrm{dup}$ denotes the quantity of duplicated episodes in OL-data phase. This table shows the BT-Fitness results.
Algorithm Batch Test Log Fitness
dup=0 dup=1 dup=2 dup=3 dup=4 dup=5 dup=6 dup=7 dup=8 dup=9
SBUCB 1 1 1 1 1 1 1 1 1 1
B-IRL 0.820±0.013 0.820±0.012 0.820±0.013 0.820±0.013 0.819±0.014 0.818±0.014 0.819±0.014 0.818±0.011 0.818±0.011 0.819±0.014
BaseSVC 0.781±0.020 0.783±0.019 0.782±0.020 0.781±0.018 0.780±0.021 0.780±0.020 0.779±0.020 0.778±0.018 0.776±0.018 0.776±0.021
IBCB (Ours) 0.963±0.029 0.964±0.015 0.960±0.027 0.967±0.017 0.967±0.013 0.972±0.015 0.975±0.014 0.953±0.034 0.973±0.012 0.973±0.013
Note that BaseSVC cannot obtain reward parameters, and thus were not included in OL-Fitness. We used ‘Train Log Behaviors Fitness’ to evaluate whether BaseSVC suffers from contradictory data and make comparison with IBCB and other baselines. As shown in Table XIII, contradictory data (same context in different time periods) undermines BaseSVC’s train log behaviors fitness, but IBCB stayed stable in OL-Fitness. That is, Contradictory data significantly affects BaseSVC’s fitness but does not impact models that learn reward parameters like IBCB, highlighting IBCB’s robustness in extreme conditions.
@@ 5.6.4 Detailed Results of Contradictory Data Testing on Synthetic Dataset
TABLE XV: Reward parameters’ estimation (Left) and training time (Right) comparison on synthetic dataset (IBCB’s parameter test). Experts do not need train since they produce train logs (behavioral evolution history logs) for baselines and IBCB. BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison. Original Expert is SBUCB.
Algorithm Online Train Log Fitness (Param Test) Algorithm Train Time(sec.) (Param Test)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 0.883±0.007 0.879±0.016 0.878±0.018 0.874±0.019 SBUCB / / / /
B-IRL 0.827±0.012 0.808±0.039 0.806±0.037 0.802±0.036 B-IRL 36.997±0.803 37.670±1.328 39.910±3.419 37.424±0.563
BaseSVC / / / / BaseSVC 116.86±20.63 116.39±26.58 132.18±19.86 117.25±19.47
B-ICB-S 0.805±0.020 0.771±0.031 0.780±0.034 0.771±0.024 B-ICB-S 86.365±0.827 87.599±4.489 86.85±1.162 85.713±0.676
IBCB-$\alpha$-1 0.868±0.008 0.858±0.036 0.859±0.038 0.861±0.029 IBCB-$\alpha$-1 0.815±0.045 0.999±0.128 0.866±0.137 1.016±0.109
IBCB-$\alpha$-0.6 0.880±0.022 0.876±0.006 0.866±0.012 0.848±0.050 IBCB-$\alpha$-0.6 0.907±0.109 0.971±0.092 0.904±0.106 0.867±0.121
IBCB-$\alpha$-0.8 0.880±0.022 0.876±0.006 0.866±0.012 0.848±0.050 IBCB-$\alpha$-0.8 0.958±0.134 0.880±0.097 0.868±0.086 0.961±0.119
IBCB-$\alpha$-1.2 0.867±0.005 0.870±0.007 0.869±0.010 0.862±0.009 IBCB-$\alpha$-1.2 0.725±0.037 0.776±0.036 0.730±0.025 0.772±0.041
IBCB-$\alpha$-1.5 0.879±0.012 0.880±0.013 0.870±0.018 0.874±0.014 IBCB-$\alpha$-1.5 0.462±0.052 0.467±0.054 0.469±0.044 0.468±0.046
IBCB-$\alpha$-2 0.865±0.007 0.865±0.012 0.864±0.016 0.861±0.008 IBCB-$\alpha$-2 0.574±0.007 0.582±0.017 0.557±0.052 0.538±0.070
TABLE XVI: Expert policy parameters’ estimation on synthetic dataset (IBCB’s parameter test). Expert policy produces same actions compared with itself, so BT-Fitness of expert policy is 1. Original Expert is SBUCB.
Algorithm Batch Test Log Fitness (Param Test) Algorithm Batch Test Average Reward (Param Test)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 1 1 1 1 SBUCB 0.641±0.010 0.629±0.028 0.630±0.024 0.622±0.024
B-IRL 0.828±0.005 0.820±0.013 0.820±0.014 0.818±0.016 B-IRL 0.588±0.015 0.571±0.037 0.572±0.034 0.561±0.034
BaseSVC 0.791±0.008 0.781±0.020 0.781±0.020 0.777±0.002 BaseSVC 0.577±0.017 0.559±0.042 0.559±0.037 0.547±0.037
B-ICB-S 0.800±0.032 0.767±0.055 0.780±0.045 0.776±0.022 B-ICB-S 0.581±0.006 0.553±0.024 0.560±0.025 0.552±0.025
IBCB-$\alpha$-1 0.970±0.012 0.960±0.036 0.959±0.034 0.969±0.020 IBCB-$\alpha$-1 0.647±0.012 0.638±0.026 0.639±0.019 0.628±0.022
IBCB-$\alpha$-0.6 0.964±0.022 0.966±0.017 0.950±0.015 0.933±0.047 IBCB-$\alpha$-0.6 0.644±0.013 0.634±0.032 0.639±0.03 0.637±0.018
IBCB-$\alpha$-0.8 0.969±0.017 0.954±0.032 0.937±0.042 0.930±0.041 IBCB-$\alpha$-0.8 0.647±0.012 0.640±0.025 0.644±0.02 0.635±0.020
IBCB-$\alpha$-1.2 0.972±0.006 0.971±0.012 0.970±0.005 0.967±0.013 IBCB-$\alpha$-1.2 0.647±0.012 0.632±0.032 0.633±0.029 0.625±0.030
IBCB-$\alpha$-1.5 0.959±0.011 0.960±0.012 0.943±0.015 0.956±0.020 IBCB-$\alpha$-1.5 0.650±0.010 0.637±0.029 0.643±0.020 0.631±0.025
IBCB-$\alpha$-2 0.970±0.006 0.973±0.017 0.962±0.018 0.960±0.021 IBCB-$\alpha$-2 0.647±0.011 0.633±0.031 0.638±0.020 0.629±0.025
5.6.5 IBCB’s Parameter Test on Synthetic Dataset
‘IBCB-$\alpha$-k’ denotes IBCB’s hyper parameter (in Eq.(17)) is k. As shown in Table XV and Table XVI, IBCB’s hyper parameter $\alpha$ will influence several indicators, especially the train time, due to the OSQP tolerance level will be relaxed if $\alpha$ is above 1.5. So for IBCB of large $\alpha$ with more relaxed tolerance, $i.e.,$ lower accuracy, training will be faster. Meanwhile, the performance of expert policy’s parameter estimation is similar according to BT-Fitness and BT-AR. By examining the parameters learned by IBCB with different $\alpha$ values, we found that after normalization (by dividing each parameter’s $l_{2}$-norm), the sum of absolute differences in every dimension of the parameters from each pair of two distinct parameters is tightly small. This suggests that one of these parameters can be a result of scaling up or down from another. During BT-data testing, items are recommended based on the parameter under BCB setting without exploration. Therefore, scaling up or down does not influence the recommended item since $\operatornamewithlimits{arg\,max}_{I\in\mathcal{I}}\langle\hat{\bm{\theta}},\bm{s}_{I}\rangle$ is only determined by the item’s content $\bm{s}_{I}$ when $\hat{\bm{\theta}}$ is fixed with a scaling factor. On the synthetic dataset, the item’s content $\bm{s}_{I}$ is very easy to separate (interval difference of each dimension of $\bm{s}_{I}$ is very obvious through the setting in Section 5.2.1), so two distinct parameters with very small sum of absolute differences may share similar BT-Fitness and BT-AR.
5.6.6 Ablation Study on ML-100K Dataset
TABLE XVII: Reward & policy parameters’ estimation on ML-100K dataset (using first half of expert’s behavioral evolution history logs). Expert policy produces same actions compared with itself, so BT-Fitness of expert policy is 1. BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison.
@@ 5.6.6 Ablation Study on ML-100K Dataset
IBCB (Ours) 0.861±0.002 0.829±0.018 0.189±0.006
TABLE XVIII: IBCB’s parameter test on ML-100K dataset. Experts do not need to train since they produce train logs (behavioral evolution history logs) for baselines and IBCB.
@@
IBCB-$\alpha$-1 0.862±0.023 0.925±0.019 0.192±0.005 4.036±0.111
IBCB-$\alpha$-0.6 0.775±0.007 0.861±0.016 0.192±0.005 3.648±0.139
IBCB-$\alpha$-0.8 0.860±0.029 0.921±0.021 0.191±0.006 4.026±0.194
IBCB-$\alpha$-1.2 0.861±0.013 0.929±0.007 0.191±0.006 4.210±0.121
IBCB-$\alpha$-1.5 0.858±0.014 0.855±0.012 0.194±0.005 2.501±0.075
IBCB-$\alpha$-2 0.845±0.017 0.855±0.012 0.194±0.005 2.663±0.221
For ML-100K dataset, Table XVII and Table XVIII reports the result of different ablation studies. Results are similar to those on synthetic dataset in most indicators: IBCB can learn from novice expert’s evolution history with better performance compared with other baselines. Note that mean and variance of train time for IBCB-$\alpha$-1.2 is larger than any other $\alpha$ settings. From the OSQP train record, we analysed that this happened due to OSQP failed to convergent for some of the OL-data train logs, so for large $\alpha$, it is recommended to use more relaxed tolerance level in OSQP setting. Note that B-ICB-S’s performance significantly decline when using first half of the train logs in Table XVII, this is because when in first half of the real-world train logs, novice expert is still exploring with large steps in a batched manner rather than updating after each step or staying stable with little exploration, which B-ICB-S assumed.
5.6.7 Extra Ablation Study Experiments of Batch Test period on ML-100K Dataset